# MaskInversion: Localized Embeddings via Optimization of Explainability Maps

### v1

@@ 4.5 Ablations
Impact of Mask Quality MaskInversion utilizes an input query mask to direct the output of the foundation model toward the area covered by the mask. Given that the mask is a critical component of our method, it is imperative to assess how variations in mask quality affect MaskInversion’s performance. To this end, we evaluate different mask conditions for the task of Class Retrieval on the MSCOCO dataset as shown in Tab. 4: Box uses the bounding boxes instead of precise segmentation masks, Box+SAM uses the bounding boxes to receive a mask via segmentation using SAM Kirillov et al. (2023), and Erosion and Dilation apply the respective morphological operations to the original masks. Fig. 5 shows qualitative examples for the different cases. Our findings indicate that eroding the mask leads to a more substantial decrease in performance compared to dilation.
We further see a decrease in accuracy from $44.7\%$ to $42.9\%$ when using bounding boxes only, whereas the combination of bounding boxes and SAM to derive the mask achieves comparable performance to the ground truth mask. This scenario is especially relevant for practical applications where users may find it easier to draw bounding boxes rather than detailed masks.
@@
Runtime Evaluation for Gradient Decomposition Tab. 5 presents a runtime comparison of the vanilla MaskInversion, where the gradient gradient-based explainability map is computed at each iteration and for each mask, versus the "gradient-decomposition" proposed in section 3.3 for $K=10$ steps. We observe that for any number of masks higher than $5$ the proposed gradient decomposition is faster than the vanilla way of computing the explainability map (see appendix Sec. E for ablation on the number of iterations).
@@ 5 Conclusion
In this work, we proposed MaskInversion as a method to create region embeddings that are grounded in the rich feature representations of foundation models without the need to fine-tune the model. To this end, we leverage the concept of explainability maps to learn an embedding vector that is focused on a respective region. We extend this idea by an add-on regularization loss to balance global and local representations as well as by a gradient decomposition to improve runtime in case of multiple masks per image. This approach holds promise for many applications in computer vision, where understanding and manipulating specific regions of an image in relation to their context is important.
@@ Acknowledgment
Walid Bousselham is supported by the German Federal Ministry of Education and Research (BMBF) project STCL-01IS22067.
Appendix A Appendix
In the Appendix, we first provide additional details on the different downstream task in Sec.B. We then provide, in Sec. C, a visualization of the explainability map throughout the optimization process. Sec.D provides visualizations of the mask distortion used for our ablations. SecE provides a more thorough ablation of the proposed gradient decomposition technique. Sec. F presents an ablation on the explainability method used for MaskInversion. Sec. G and H respectively discuss the limitations of SOTA methods as well as the proposed MaskInversion. Eventually, we provide additional qualitative examples of localized captioning and diffusion in Sec. I and Sec. J.
@@ Appendix B Downstream Tasks
To assess the proposed method’s ability to capture localized properties, we evaluate it for referring expression classification. Given an image and a set of masks, we generate an embedding for each mask within an image and match the generated region embeddings to a set of text queries (referring expressions) encoded with the respective text encoder. The query mask whose localized embedding exhibits the highest cosine similarity with the text embedding is selected. We employ standard referring expression datasets, i.e. PhraseCut Wu et al. (2020), RefCOCO, and RefCOCO+ Kazemzadeh et al. (2014). For RefCOCO and RefCOCO+, we use the mask annotations from the MSCOCO Lin et al. (2014) dataset, which has about 30 masks per image, thereby increasing the difficulty of the task. For PhraseCut, we consider the masks of all annotated referring expressions as candidates, reporting top-1, top-5, and top-10 accuracy. Additionally, following (Subramanian et al., 2022; Sun et al., 2024; Yang et al., 2024; Shtedritski et al., 2023), for RefCOCO and RefCOCO+, we report the mean Intersection over Union (mIoU) and overall Intersection over Union (oIoU).
@@
Second, we consider the task of zero-shot classification as a common benchmark for vision-language models. In that task, an image is classified by matching its visual embedding with the textual description of the classes present in the dataset. Here, we propose to increase the granularity by using it to classify a specific region of the image: given a query mask of an object, classify it by matching its localized embedding to the text embeddings of the classes in the datasets. For this, we leverage two semantic segmentation datasets, PascalVOC Everingham et al. (2015) and PascalContext Mottaghi et al. (2014), with 19 and 59 classes, respectively, and one instance segmentation dataset, MSCOCO Lin et al. (2014), with 80 classes. The performance is evaluated using the top-1, top-5, and top-10 accuracy metrics, denoted by $Acc@1$, $Acc@5$, and $Acc@10$. Finally, we challenge the proposed method in a large-scale open-vocabulary setting by using a dataset encompassing a substantially larger number of classes. We utilize a subset of the OpenImagesV7 Benenson & Ferrari (2022) dataset, which offers mask annotations for a diverse array of objects across 350 unique classes. The evaluation metrics are again top-1, top-5, and top-10 accuracy reported as $Acc@1$, $Acc@5$, and $Acc@10$.
@@
Traditionally, image captioning models generate captions for entire images based on the visual representation provided by an image encoder. In contrast, we aim to evaluate our method’s ability to focus the captioner on a specific image region while maintaining contextual relevance. To this end, we leverage a pretrained image captioner, CLIPCap Mokady et al. (2021), and provide it with the localized embedding token of a query mask to generate a caption. CLIPCap is trained on top of the CLIP vision encoder and feeds its [CLS] token to GPT-2Radford et al. (2019) to produce a caption. Here, we feed the localized embeddings of MaskInversion as a drop-in replacement of the CLIP [CLS] token to the captioner without any finetuning. As no dataset directly supports this evaluation type, we adapted an existing dataset, PhraseCut. To quantitatively evaluate the generated localized captions, we match the generated caption to the set of ground truth referring expressions for this image using the text encoder from CLIP (ViT-L/14 by OpenAI), consider the caption correct if the cosine similarity between the generated caption and the ground truth referring expression for this mask is the highest. The reported metric for this task is the top-1 accuracy.
@@ Appendix C Optimization steps visualization
Figure 4 provides a visualization of the explainability map throughout the optimization process employed by MaskInversion. It is observed that the explainability map increasingly concentrates on the region covered by the query mask as the optimization progresses. This observation is indicative of the method’s ability to effectively focus the attention of the underlying foundation model on the designated areas of the image.
Appendix D Mask Quality
@@
Appendix E Gradient Decomposition
@@
Appendix F Impact of the Explainability Method
Given that MaskInversion leverages an explainability method to guide the inversion process, its dependency on the choice of explainability method was evaluated. We experimented with alternative gradient-based methods, such as GradCAM and CheferCAM, in place of the originally used LeGrad. The comparative results on the MSCOCO dataset are presented in Table 6. LeGrad significantly outperformed the other methods, which can be attributed to its design specificity for ViT architectures, unlike GradCAM and CheferCAM, which are tailored for CNNs and general transformers, respectively. This finding aligns with the observations in Bousselham et al. (2024), where LeGrad demonstrated superior localization capabilities essential for the tasks addressed by MaskInversion. Thus, the selection of an appropriate explainability method is crucial for optimizing the performance of MaskInversion.
Appendix G SOTA Methods’ Limitations
@@
FGVPYang et al. (2024) Heavily blur the background, thus losing the context.
RISYu et al. (2023) Masks the features of the ViT after a certain number of layers to prevent
@@
AlphaCLIPSun et al. (2024) Finetunes CLIP to take as input an image and a mask.
@@ Appendix H Limitations
Appendix I Additional Localized Captions
@@
Appendix J Additional Localized Diffusion

### v2

@@ 4.4 Ablations
Impact of Mask Quality MaskInversion utilizes an input query mask to direct the output of the foundation model toward the area covered by the mask. Given that the mask is a critical element of the MaskInversion method, we explore here how variations in mask quality affect its performance. To this end, we evaluate different mask conditions for the task of Class Retrieval on the MSCOCO dataset as shown in Table 6 and Figure 9: Box uses the masks’ bounding-boxes instead of precise segmentation masks, Box+SAM feeds the bounding-boxes to SAM (Kirillov et al., 2023) to produce approximate masks and uses those instead of ground-truth masks, and Erosion and Dilation apply the respective morphological operations to the original masks. The results indicate that eroding the mask leads to a more substantial decrease in performance compared to dilation. We further see a decrease in accuracy from $44.7\%$ to $42.9\%$ when using bounding-boxes only, whereas the combination of bounding-boxes and SAM to derive the mask achieves comparable performance to inputting ground-truth mask. This scenario is especially relevant for practical applications where users may find it easier to draw bounding-boxes rather than detailed masks.
@@
Runtime Evaluation for Gradient Decomposition Table 6 presents a runtime comparison of the vanilla MaskInversion, where the gradient gradient-based explainability map is computed at each iteration for each mask, versus the ”gradient-decomposition” proposed in section 3.2 for $K=10$ steps. When there are more than 5 masks in an image, the proposed gradient decomposition is faster than the vanilla way of computing the explainability map (see appendix Sec. I for an ablation on the number of iterations).
4.5 Localized Captioning Analysis
We further consider the performance of MaskInversion against CLIP and AlphaCLIP for localized captioning in Table 6. We start from CLIPCap (Mokady et al., 2021) as the base captioner and replace the CLIP image encoder with either AlphaCLIP or the output of MaskInversion without any fine-tuning. We observe that MaskInversion demonstrates the ability to focus the captioner on the region of interest, as the accuracy more than doubles when using MaskInversion versus only using CLIP. Moreover, MaskInversion also significantly outperforms AlphaCLIP, despite not involving any fine-tuning of the CLIP model.
Figure 3: Localized Embedding Visualizations: Visualisation of the learned localized embedding using (left) a pretrained diffusion model; (right) an image captioner. In both cases, the global feature representation is replaced by the output of MaskInversion depending on the query mask.
Figure 3 presents qualitative examples of the localized captions generated by MaskInversion+CLIPCap for different query masks. MaskInversion demonstrates a high degree of precision in focusing the captioning model on specific image regions specified by the query masks. Both, the caption and the heatmap, focus on the area covered by the query mask.
4.6 Mask Embedding for Image Diffusion
To further visualize the concepts captured in the learned representation output by MaskInversion, we employed $\lambda$-ECLIPSE (Patel et al., 2024), which takes as input a visual embedding from a ViT-bigG/14 CLIP model along with a text prompt and generates variations of the input image that correspond to the prompt. Utilizing the default settings of $\lambda$-ECLIPSE as described in (Patel et al., 2024), we generate images based on different query masks used within the MaskInversion process. As Figure 3 show, resulting images vary depending on the mask used. The images focus on the objects inside the query mask, confirming that MaskInversion directs the model’s attention to specific parts of the image. Moreover, we observe that the final explainability map is focused on the area covered by the mask, validating the effectiveness of our proposed optimization process.
@@ 5 Conclusion
We proposed MaskInversion as a method to create region embeddings that are grounded in the rich feature representations of foundation models, without the need to fine-tune the model. To this end, we leveraged explainability maps to learn an embedding vector focused on a specific image region. We extend this idea with an add-on regularization loss to balance global and local representations, and with a gradient decomposition technique to improve runtime.
@@ Acknowledgment
Walid Bousselham is supported by the German Federal Ministry of Education and Research (BMBF) project STCL - 01IS22067. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer JUPITER — JUWELSJülich Supercomputing Centre (2021) at Jülich Supercomputing Centre (JSC). This work also acknowledges support from the project ELLIOT-FM: Open Multi-Modal Foundation Models with Strong Generalization and Reasoning.
Appendix
Appendix A Overview
@@
In the Appendix, we first provide additional details on the different downstream tasks in Sec.B. Sec.C provides a visualization of the explainability map throughout the optimization process. Sec.4.3 presents a comprehensive comparison with training-free methods. Sec.F analyzes the influence of the hyperparameter $\alpha$ on balancing local and global information. Sec.G demonstrates our method’s capability to handle multiple objects. Sec.H provides visualizations of the mask distortion used for our ablations. Sec.I presents an ablation of the proposed gradient decomposition technique. Sec.J evaluates the impact of different explainability methods on MaskInversion. Sec.K and Sec.L respectively discuss the limitations of SOTA methods and the proposed MaskInversion. Finally, we provide additional qualitative examples of localized captioning and diffusion in Sec.M and Sec.N.
@@ Appendix B Downstream Tasks
To assess the proposed method’s ability to capture localized properties, we evaluate it for referring expression classification. Given an image and a set of masks, we generate an embedding for each mask within an image and match the generated region embeddings to a set of text queries (referring expressions) encoded with the respective text encoder. The query mask whose localized embedding exhibits the highest cosine similarity with the text embedding is selected. We employ standard referring expression datasets, i.e. PhraseCut (Wu et al., 2020), RefCOCO, and RefCOCO+ (Kazemzadeh et al., 2014). For RefCOCO and RefCOCO+, we use the mask annotations from the MSCOCO (Lin et al., 2014) dataset, which has about 30 masks per image, thereby increasing the difficulty of the task. For PhraseCut, we consider the masks of all annotated referring expressions as candidates, reporting top-1, top-5, and top-10 accuracy. Additionally, following (Subramanian et al., 2022; Sun et al., 2024; Yang et al., 2023; Shtedritski et al., 2023), for RefCOCO and RefCOCO+, we report the mean Intersection over Union (mIoU) and overall Intersection over Union (oIoU).
@@
Second, we consider the task of zero-shot classification as a common benchmark for vision-language models. In that task, an image is classified by matching its visual embedding with the textual description of the classes present in the dataset. Here, we propose to increase the granularity by using it to classify a specific region of the image: given a query mask of an object, classify it by matching its localized embedding to the text embeddings of the classes in the datasets. For this, we leverage two semantic segmentation datasets, PascalVOC (Everingham et al., 2015) and PascalContext (Mottaghi et al., 2014), with 19 and 59 classes, respectively, and one instance segmentation dataset, MSCOCO (Lin et al., 2014), with 80 classes. The performance is evaluated using the top-1, top-5, and top-10 accuracy metrics, denoted by $Acc@1$, $Acc@5$, and $Acc@10$. Finally, we challenge the proposed method in a large-scale open-vocabulary setting by using a dataset encompassing a substantially larger number of classes. We utilize a subset of the OpenImagesV7 (Benenson and Ferrari, 2022) dataset, which offers mask annotations for a diverse array of objects across 350 unique classes. The evaluation metrics are again top-1, top-5, and top-10 accuracy reported as $Acc@1$, $Acc@5$, and $Acc@10$.
@@
Traditionally, image captioning models generate captions for entire images based on the visual representation provided by an image encoder. In contrast, we aim to evaluate our method’s ability to focus the captioner on a specific image region while maintaining contextual relevance. To this end, we leverage a pretrained image captioner, CLIPCap (Mokady et al., 2021), and provide it with the localized embedding token of a query mask to generate a caption. CLIPCap is trained on top of the CLIP vision encoder and feeds its [CLS] token to GPT-2(Radford et al., 2019) to produce a caption. Here, we feed the localized embeddings of MaskInversion as a drop-in replacement of the CLIP [CLS] token to the captioner without any finetuning. As no dataset directly supports this evaluation type, we adapted an existing dataset, PhraseCut. To quantitatively evaluate the generated localized captions, we match the generated caption to the set of ground truth referring expressions for this image using the text encoder from CLIP (ViT-L/14 by OpenAI), consider the caption correct if the cosine similarity between the generated caption and the ground truth referring expression for this mask is the highest. The reported metric for this task is the top-1 accuracy.
@@ Appendix C Optimization steps visualization
Finally, Figure 4 provides a visualization of the explainability map throughout the optimization process employed by MaskInversion. It is observed that the explainability map increasingly concentrates on the region covered by the query mask as the optimization progresses. This observation is indicative of the method’s ability to effectively focus the attention of the underlying foundation model on the designated areas of the image.
Appendix D Convergence Analysis and Hyperparameter Choice
In Section 3.2, we set the number of gradient descent iterations for the optimization of the localized embedding token at $K=10$. This value was determined through an empirical analysis designed to balance the quality of the learned embedding with computational efficiency.
To illustrate this choice, we monitor the optimization process on the PascalVOC dataset. Figure 5 presents the evolution of both the optimization loss (left y-axis, red curve) and the downstream classification accuracy (right y-axis, blue curve) over 100 iterations.
Figure 5: Convergence Analysis of MaskInversion. The plot illustrates the optimization loss (red, left axis) and the resulting accuracy on PascalVOC (blue, right axis) over iterations. The dotted line marks the chosen stopping point at $K=10$ iterations.
As demonstrated in Figure 5, the optimization process converges quickly. The classification accuracy increases sharply in the initial steps, reaching approximately $85\%$ by iteration 10, after which it effectively plateaus. This asymptotic behavior indicates that the localized embedding token $LET_{\mathbf{m}}$ quickly captures the core semantic features required for the task, with the explainability map becoming sufficiently aligned with the query mask within the first 10 iterations.
While the loss function continues to decrease marginally beyond the $10^{\text{th}}$ iteration, these residual improvements primarily correspond to boundary refinements between the generated explainability map and the binary query mask. As evidenced by our ablation experiments on mask quality (see Table 6 in the main text), such minor boundary adjustments have a minimal impact on the semantic quality of the learned embedding. Consequently, extending the optimization beyond this point incurs additional computational cost without yielding significant performance gains. Therefore, we select $K=10$ to ensure high-quality localized embeddings while maintaining computational efficienty.
Appendix E Empirical Runtime Analysis
A key consideration for high-resolution tasks, such as medical imaging or aerial photography, is how the computational cost of this method scales with increased input resolution. To empirically validate this scalability, we measured the inference runtime while varying the input image resolution. We tested standard resolutions of $224\times 224$ pixels and increased up to $768\times 768$ pixels. We a patch size of $16\times 16$, this corresponds to a range of $n=196$ to $n=2,304$ visual tokens. The experiment was conducted with a fixed budget of $K=10$ iterations processing 10 masks per image.
Figure 6: Runtime Scalability. The plot displays the processing time (in seconds) as a function of the number of visual tokens $n$. The input resolutions correspond to $224^{2}$, $448^{2}$, $512^{2}$, $640^{2}$, and $768^{2}$. The results demonstrate a clear linear relationship between runtime and the number of tokens.
The results are illustrated in Figure 6. As predicted, the runtime shows a linear scaling with the number of visual tokens. For instance, increasing the resolution such that the token count doubles results in an approximate doubling of the processing time. This linear scaling confirms that MaskInversion remains computationally tractable even for higher-resolution inputs.
Appendix F Influence of $\alpha$
Figure 7: Qualitative analysis of the influence of $\alpha$ on the generated captions.
alpha 0.0 0.5 1.0 1.5 2.0 2.5 3.0 4.0 5.0 5.5 6.0 6.5 7.0 7.5 8.0 10.0 20.0
Acc 41.7 47.6 50.3 52.2 53.9 54.6 55.2 56.0 56.2 56.0 55.8 56.0 56.2 56.1 55.8 53.7 20.5
Table 7: Accuracy for different values of $\alpha$ on RefCOCO.
We conduct an extensive analysis of the hyperparameter $\alpha$ to understand its role in balancing local and global information within the learned embeddings. Figure 7 illustrates this effect through generated captions for different $\alpha$ values. When $\alpha=0$, the model generates descriptions focused strictly on the masked region (e.g., “woman in a boat”), while increasing $\alpha$ progressively incorporates more contextual information(e.g., “produce” or “vegetables”). Quantitatively, we observe that performance on RefCOCO improves as $\alpha$ increases from 0 (41.7%) to an optimal value around $\alpha=5.0$ (56.2%), before gradually declining for larger values. This sweet spot ($\alpha\approx 5.0$) represents an optimal balance where the embedding retains sufficient local information while leveraging beneficial contextual cues. Beyond $\alpha>7.5$, performance deteriorates as the representation becomes increasingly similar to the global [CLS] token, with a dramatic drop at $\alpha=20.0$ (20.5%). This analysis demonstrates that $\alpha$ effectively functions as a control mechanism for trading off local detail against global context in the learned representations.
Appendix G Multi-Object
Figure 8: Multi-Object Analysis: Visualization of MaskInversion’s ability to handle multiple objects. (top) Query masks highlighting different combinations of objects, (middle) corresponding heatmaps showing the model’s focus regions, and (bottom) generated images using $\lambda$-ECLIPSE demonstrating the preservation of multiple object characteristics in the learned embeddings.
While quantitative evaluation of multi-object scenarios presents inherent challenges, we demonstrate MaskInversion’s capability to handle multiple objects through qualitative analysis. As shown in Figure 8, our method effectively captures the relationships and context of multiple objects within a single mask. For instance, when given a mask covering multiple Pokémon characters, the generated diffusion outputs maintain coherent representations of all objects while preserving their spatial relationships and individual characteristics. The diffusion model successfully reconstructs multiple objects from the localized embedding, indicating that MaskInversion effectively encodes information about multiple entities and their relative positioning. This is particularly evident in cases where the mask encompasses groups of similar objects (e.g., multiple Pokémon) or diverse object combinations, demonstrating the method’s robustness in handling complex, multi-object scenarios without losing individual object details or their contextual relationships .
Appendix H Mask Quality
@@
Appendix I Gradient Decomposition
@@
I.1 Detailed Derivation of Gradient Decomposition
In Section 3.2, we introduced a gradient decomposition strategy to enhance the computational efficiency of the MaskInversion process. This strategy relies on the independence of the localized embedding token $LET_{\mathbf{m}}^{(k)}$ from the model activations $\mathbf{A}$ during the gradient computation step. Here, we provide a more detailed formal proof supporting Equation 5.
Problem Formulation
Let the scalar activation score $s$ be defined as the dot product between the global image representation $\bar{\mathbf{z}}$ and the localized embedding token $LET_{\mathbf{m}}^{(k)}$ at iteration $k$. For clarity, we omit normalization terms (as in cosine similarity), noting that this does not alter the dependency logic:
$s=\bar{\mathbf{z}}\cdot(LET_{\mathbf{m}}^{(k)})^{T},$ (6)
where:
  • $\bar{\mathbf{z}}=\frac{1}{n}\sum_{p}z_{p}:=f(\mathbf{A})$ represents the combined patch and [CLS] token representation averaged across spatial dimensions. This term is a direct function of the intermediate activations $\mathbf{A}$.
  • $LET_{\mathbf{m}}^{(k)}$ represents the learnable vector parameters at the current optimization step $k$.
Derivation
Our objective is to compute the gradient of the score $s$ with respect to the activations $\mathbf{A}$, denoted as $\nabla\mathbf{A}$. By applying the product rule of calculus, the gradient can be expanded as follows:
$\nabla\mathbf{A}=\frac{\partial s}{\partial\mathbf{A}}=\frac{\partial\left(\bar{\mathbf{z}}\cdot(LET_{\mathbf{m}}^{(k)})^{T}\right)}{\partial\mathbf{A}}$ (7)
Expanding the terms yields:
$\frac{\partial s}{\partial\mathbf{A}}=\left(\frac{\partial\bar{\mathbf{z}}}{\partial\mathbf{A}}\cdot(LET_{\mathbf{m}}^{(k)})^{T}\right)+\left(\bar{\mathbf{z}}\cdot\frac{\partial(LET_{\mathbf{m}}^{(k)})^{T}}{\partial\mathbf{A}}\right).$ (8)
Proof of Independence
To evaluate the second term in Equation equation 8, we examine the definition of the localized token. While $LET_{\mathbf{m}}^{(0)}$ may be initialized using the global [CLS] token (which is derived from $\mathbf{A}$), in our implementation, the token is strictly detached from the computational graph upon initialization:
$LET_{\mathbf{m}}^{(0)}:=\mathrm{stop\_gradient}(\text{[CLS]}).$ (9)
For all subsequent iterations $k$, $LET_{\mathbf{m}}^{(k)}$ is treated as an external, standalone optimization variable updated via the optimizer, rather than a continuous function of the image input from the current forward pass. Consequently, the partial derivative of the token with respect to the current activations is zero:
$\frac{\partial LET_{\mathbf{m}}^{(k)}}{\partial\mathbf{A}}=\mathbf{0}.$ (10)
Substituting this into Equation equation 8, the second term vanishes:
$\nabla\mathbf{A}$ $=\left(\frac{\partial\bar{\mathbf{z}}}{\partial\mathbf{A}}\cdot(LET_{\mathbf{m}}^{(k)})^{T}\right)+\left(\bar{\mathbf{z}}\cdot\mathbf{0}\right)$ (11)
$=\frac{\partial\bar{\mathbf{z}}}{\partial\mathbf{A}}\cdot(LET_{\mathbf{m}}^{(k)})^{T}.$
This derivation confirms that the decomposition presented in Equation (5) is mathematically exact. It allows for the pre-computation of the Jacobian $\frac{\partial\bar{\mathbf{z}}}{\partial\mathbf{A}}$, which can then be reused across all $K$ iterations via a simple dot product with the evolving token $LET_{\mathbf{m}}^{(k)}$, reducing computational cost.
Appendix J Impact of the Explainability Method
@@ I.1 Detailed Derivation of Gradient Decomposition
Given that MaskInversion leverages an explainability method to guide the inversion process, its dependency on the choice of explainability method was evaluated. We experimented with alternative gradient-based methods, such as GradCAM and CheferCAM, in place of the originally used LeGrad. The comparative results on the MSCOCO dataset are presented in Table 8. LeGrad significantly outperformed the other methods, which can be attributed to its design specificity for ViT architectures, unlike GradCAM and CheferCAM, which are tailored for CNNs and general transformers, respectively. This finding aligns with the observations in (Bousselham et al., 2025), where LeGrad demonstrated superior localization capabilities essential for the tasks addressed by MaskInversion. Thus, the selection of an appropriate explainability method is crucial for optimizing the performance of MaskInversion.
Appendix K SOTA Methods’ Limitations
@@
FGVP(Yang et al., 2023) Heavily blur the background, thus losing the context.
RIS(Yu et al., 2023) Masks the features of the ViT after a certain number of layers to prevent
@@
AlphaCLIP(Sun et al., 2024) Finetunes CLIP to take as input an image and a mask.
@@
Appendix M Additional Localized Captions
@@
Appendix N Additional Localized Diffusion